OpenStudy (anonymous):

can someone check my work on my previous post about the 420 pounds? can someone check my work on my previous post about the 420 pounds? @Mathematics

OpenStudy (turingtest):

can you post the link?

OpenStudy (anonymous):

if you scroll down some in the room its still showing as posted :)

OpenStudy (anonymous):

i will post it here

OpenStudy (anonymous):

an object is 420 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.) can someone help me work this out please I dont understand any of it

OpenStudy (anonymous):

ive got that part figured out I think its just the rest of it

OpenStudy (turingtest):

It seems we are missing some context here. What kind of problems have you been doing in class? Physics, linear equations, proportions, etc?

OpenStudy (anonymous):

we are doing like the square root, im new to it all

OpenStudy (anonymous):

w=Cr^-2

OpenStudy (turingtest):

Ahhhhhh the equation!

OpenStudy (turingtest):

that makes this doable.

OpenStudy (anonymous):

so I have w=420(3963)^-2

OpenStudy (turingtest):

your equation is w=Cr^-2 now which variable represents which number?

OpenStudy (anonymous):

w=420(15705369) w=6596254980

OpenStudy (turingtest):

I mean w=? r=? C=? what does each variable represent?

OpenStudy (anonymous):

the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth.

OpenStudy (anonymous):

so r would be 3963?

OpenStudy (turingtest):

I'm sorry, I have no idea how you got that number. As to your previous comment you are right. So then which number do we NOT know and need to find out

OpenStudy (turingtest):

?

OpenStudy (turingtest):

That is, which numbers do we know, and which do we not know. Do we know the value of r? Do we know the value of C?

OpenStudy (anonymous):

maybe i have confused you lol let me do it like this... the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth. Solve the equation for r. . Suppose that an object is 420 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.)

OpenStudy (anonymous):

1 mile = 5,280 feet

OpenStudy (anonymous):

can you see why i have no clue what im doing lol

OpenStudy (turingtest):

why do you think you need to solve for r? In math we solve for the variable we do NOT know the value of, in this case C. We need to find what value of C makes the statement above true, which is done by solving for C: w=Cr^-2 C=wr^2 now plug in the values for w and r and this equation will give you C

OpenStudy (anonymous):

oh so i needed to multiply or something to get it in that form?

OpenStudy (anonymous):

c=420(3963)^2

OpenStudy (turingtest):

yes, the operation to go from w=Cr^-2--->C=wr^2 is that we multiplied both sides by r^2 w(r^2)=Cr^(-2)(r^2) wr^2=C if you understand adding exponents this should make sense yes C=420(3963)^2=659625980 I am not sure of the units, but that is the number

OpenStudy (anonymous):

. Now i Use the value of C i found in the previous question to determine how much the object would weigh In Moscow, Russia (450 feet above sea level).

OpenStudy (turingtest):

did you give up?

OpenStudy (anonymous):

soo 65962598(3963-450/5280)^2

OpenStudy (anonymous):

no im still working on it lol

OpenStudy (turingtest):

ok cool

OpenStudy (turingtest):

I can give hints if you want

OpenStudy (anonymous):

65962598(3962.9148)^2 65962598(15704655.6681

OpenStudy (anonymous):

if you can just see if im doing it right and just help me work it out would be fine

OpenStudy (anonymous):

i just dont think what im doing is right though

OpenStudy (turingtest):

I'm sorry but you're quite far off. Our equation is \[w=Cr^{-2}={C \over r^2}\] we know r in moscow (it's r at sea level plus 450) and after our last problem we know C so which numbers are we going to plug in?

OpenStudy (anonymous):

450 and 3963?

OpenStudy (anonymous):

no thats not right

OpenStudy (turingtest):

correct so where and what is the answer?

OpenStudy (anonymous):

oh it is right lol

OpenStudy (anonymous):

w=450(3963)^2?

OpenStudy (turingtest):

I', sorr, those number are correct for the new r, (that is r of moscow) in moscow, r=3963+450 AND we need to plug in C to the right place. My fault gor not mentioning that. So what is the answer?

OpenStudy (anonymous):

4413

OpenStudy (anonymous):

dont i have to divide the 450 into the 5280?

OpenStudy (turingtest):

im sorry my connection is bad and this server is as well one moment, it's on paper but hard to type

OpenStudy (anonymous):

what if i emailed you the sheet so you can look at it ans see what im talking about. I feel like ive confused you a ton lol

OpenStudy (anonymous):

wait i can do it here i think

OpenStudy (turingtest):

\[w={C \over r^2}\]\[C=w r^2=420(3963)^2\] for moscow \[w={C \over r_m2^2}={420*3963^2\over(3963+450)^2}=338.2\]I'm sorry this answer must be wrong and makes no sense. It's very late where I am and I'm more tired than I thought. I'll help you next time I promise.

OpenStudy (turingtest):

): goodnight...

OpenStudy (anonymous):

goodnight

OpenStudy (anonymous):

thanks for the help though